OVERVIEW OF DOE COMBUSTION PROGRAM \underline{WILLIAM H. KIRCHHOFF}, Chemical Sciences Division, Department of Energy, ER-141 GTN MS G-23, Washington D.C. 20585.

This talk will breifly outline the mission and goals of the DOE combustion program. Particular attention will be devoted to the relevance of modern spectroscopic research to the program.